how many liters of a 40% alcohol solution must be mixed 60 liters of a 70% solution to get a 50% solution
Answer by stanbon(75887)

how many liters of a 40% alcohol solution must be mixed 60 liters of a 70% solution to get a 50% solution
-------
Using 2 equation:
Quantity Eq:: x + 60 = y
Alcohol Eq:::0.40x + 0.70*60 = 0.50y
------------
Multiply thru Quantity Eq by 40
Multiply thru Alcohol Eq by 100
------
40x + 40*60 = 40y
40x + 70*60 = 50y
-----
Subtract and solve for "y":
30*60 = 10y
y = 180 liters
------
Solve for "x":
x + 60 = y
x + 60 = 180
x = 120 liters (amt. of 40% solution needed)
======================
Cheers,
Stan H.